# Catastrophic Cyber Capabilities Benchmark (3CB): Robustly Evaluating LLM Agent Cyber Offense Capabilities

### v1

@@ 1 Introduction
Artificial intelligence (AI), particularly large language models (LLMs), is rapidly advancing in capabilities (Epoch AI, 2023). AI is integrated into applications, including chatbots, coding assistants, and autonomous programming agents (Wang et al., 2024). While these models offer substantial benefits and have the potential to revolutionize industries, their dual-use nature—capable of being leveraged for both beneficial and harmful purposes—raises significant concerns (Hendrycks et al., 2024). The potential misuse of AI in cyber offense operations is increasingly alarming, with 93% of cybersecurity experts predicting AI-induced ”cyber catastrophes” by 2026, such as critical infrastructure breakdown and ransomware (Forum, 2023). Moreover, a survey of over 2,000 AI researchers revealed that 41% anticipate that human-level machine intelligence could pose existential risks to humanity within the next century (Grace et al., 2024). These projections underscore the need for comprehensive evaluations of AI capabilities to mitigate potential risks and ensure safer deployment.
@@
  • We introduce 3CB, the first cyber offense capability benchmark designed to represent all skills relevant to the cybersecurity domain, with 15 original challenges (Section 2.3).
@@
  • We find that cyber offense performance is highly variable and conditional on subtle changes to prompting and the environment (Figure 5).
@@ 1.1 Related Work
Phuong et al. (2024) develop a series of capture-the-flag (CTF) challenges representing realistic scenarios involving web application vulnerabilities and privilege escalation. These efforts often lack a structured taxonomy, limiting the benchmark’s generalization. Such collections are also prone to not being representative, as there is a risk of models having memorized public CTF challenges Fang et al. (2024) demonstrates the alarming capabilities of LLMs in cyber offense. They show that tool-augmented LLMs can autonomously exploit vulnerabilities in sandboxed websites. Their results highlight the urgent need for robust evaluations; GPT-4 is able to hack 73% of the websites with its predecessor of barely a year, GPT-3.5, reaching just 7%. All open source models fail at this task. Supporting this, Pa Pa et al. (2023) find that current LLM services are not properly safeguarded against supporting malware attacks.
@@ 1.2 Scenarios for Catastrophic AI Cyber Risk
The intersection of catastrophic AI risk and cybersecurity has gained significant attention, both academically (Grace et al., 2024) and publicly (Forum, 2023). Cyber offense capabilities have led to critical infrastructure disruptions, such as power outages in Ukraine (Whitehead et al., 2017), more than \$1.8 billion damages from a single malware attack (Crosignani et al., 2024), and intellectual property theft causing hundreds of billions in damages in the US alone (FBI, 2019). The increasing reliance on digital infrastructure, combined with the proliferation of cyber threats and potential AI risks, presents a formidable challenge. Recognizing these emerging threats, various stakeholders are taking proactive measures:
  • Research Laboratories and Academia: Developing classification systems and risk assessment methodologies for foundation models’ cyber capabilities to enable preemptive security interventions (Li et al., 2024; Phuong et al., 2024; Bhatt et al., 2024; OpenAI, 2024).
  • Governments: Investing in foundation model evaluations and AI safety research, focusing on high-risk areas like cyber offense and chemical/biological capabilities (DSIT, 2024; Nist, 2024).
  • AI Development Companies: Implementing security and risk mitigation strategies based on risk evaluations, including policies that establish safeguards at specific risk levels (Anthropic, 2023; OpenAI, 2023; Morris et al., 2024).
Despite these efforts, there is a lack of a comprehensive, structured, principled evaluation of LLMs’ general cyber offensive capabilities. So we created 3CB.
@@ 6 Ethics Statement
We acknowledge that our work is creating agents and infrastructure that might be misused. We decide to release these due to their limited performance gap compared to raw LLM queries. Due to the potential for threat actors to use our scaffolding and challenges to train frontier agents, we avoid releasing the four most difficult challenges: sshhijack, bashhist, nodecontrol and rce.

### v2

@@ 1 Introduction
Artificial intelligence (AI), particularly large language models (LLMs), is rapidly advancing in capabilities (Epoch AI, 2023; OpenAI et al., 2024; Anthropic, 2024a). AI is integrated into applications, including chatbots, coding assistants, and autonomous programming agents (Wang et al., 2024). While these models offer substantial benefits and have the potential to revolutionize industries, their dual-use nature—capable of being leveraged for both beneficial and harmful purposes—raises significant concerns (Hendrycks et al., 2024). The potential misuse of AI in cyber offense operations is increasingly alarming, with 93% of cybersecurity experts predicting AI-induced ”cyber catastrophes” by 2026, such as critical infrastructure breakdown and ransomware (Forum, 2023). Moreover, a survey of over 2,000 AI researchers revealed that 41% anticipate that human-level machine intelligence could pose existential risks to humanity within the next century (Grace et al., 2024). These projections underscore the need for comprehensive evaluations of AI offense capabilities to mitigate potential risks and ensure safer deployment.
@@
  • We introduce 3CB, the first cyber offense capability benchmark designed to represent all skills relevant to the cyber offense domain, with 15 original challenges (Section 2.3).
@@
  • We find that cyber offense performance is highly variable and conditional on subtle changes to prompting and environment variations (Figure 5).
@@ 1.1 Related Work
Interactive cybersecurity challenge environments for LLMs: Phuong et al. (2024) develop a series of capture-the-flag (CTF) challenges representing realistic scenarios involving web application vulnerabilities and privilege escalation. Fang et al. (2024) show that tool-augmented LLMs can autonomously exploit vulnerabilities in sandboxed websites. GPT-4 is able to hack 73% of the websites with its predecessor of barely a year, GPT-3.5, reaching just 7%. All open source models fail at this task. Zhang et al. (2024) find LLMs perform well on capture-the-flag (CTF) challenges from four competitions and perform fine-grained evaluation using sub-tasks for each task. Yang et al. (2023) create interactive environments for LLMs in bash, SQL and Python to evaluate model performance on related tasks in each setting. Shao et al. (2024) compile an open repository of CTF challenges from online sources and create an interactive CTF playground for models.
Pa Pa et al. (2023) find that current LLM services are not properly safeguarded against cyber offense misuse. Haimes et al. (2024) show that publicly accessible benchmark content may be memorized by models, leading to untrustworthy performance on benchmarks.
@@ 1.2 Scenarios for Catastrophic AI Cyber Risk
The intersection of catastrophic AI risk and cybersecurity has gained significant attention, both academically (Grace et al., 2024) and publicly (Forum, 2023). Cyber offense capabilities have led to critical infrastructure disruptions, such as power outages in Ukraine (Whitehead et al., 2017), more than \$1.8 billion damages from a single malware attack (Crosignani et al., 2024), and intellectual property theft causing hundreds of billions in damages in the US alone (FBI, 2019). In the face of increasing reliance on digital infrastructure and heightening cyber crime:
  • Research laboratories and academia are developing classification systems and risk assessment methodologies for foundation models’ cyber capabilities to enable preemptive security interventions (Li et al., 2024; Phuong et al., 2024; Bhatt et al., 2024; OpenAI, 2024).
  • Governments are investing in foundation model evaluations and AI safety research, focusing on high-risk areas like cyber offense and chemical/biological capabilities (DSIT, 2024; Nist, 2024).
  • AI development companies are implementing security and risk mitigation strategies based on risk evaluations, including policies that establish safeguards at specific risk levels (Anthropic, 2023; OpenAI, 2023; Morris et al., 2024).
Despite these efforts, there is a lack of a memorization-safe and principled evaluation of LLMs’ general cyber offensive capabilities. So we created 3CB.
@@ 6 Ethics Statement
We acknowledge that our work introduces agents and infrastructure that is at risk of being misused. We decide to release these due to their limited performance gap compared to raw LLM queries. Due to the potential for threat actors to use our scaffolding and challenges to train frontier agents, we avoid releasing the four most difficult challenges: sshhijack, bashhist, nodecontrol and rce.